This question is from my beginning algebra textbook and I have tried it two different ways and still can't seem to come up with the right answer.
Here is the problem: In January 2006 U.S. first-class mail rates increased to 39 cents for the first ounce, plus 24 cents for each additional ounce. If Sabrina spent $15.00 for a total of 45 stamps of these two denominations, how many stamps of each denomination did she buy?
I used x for the 39 cent stamps and x-15 for the 24 cent stamps and have tried to work it out like this: x+x-15+45, that didn't work so I tried it this way: x+x-15=$15.00 and that didn't work either. I know the answer is 28 39 cent stamps and 17 24 cent stamps but I like to know how to work them out because I learn better that way. Please help me to work it out the right way.

Answer by stanbon(75887)

This question is from my beginning algebra textbook and I have tried it two different ways and still can't seem to come up with the right answer.
Here is the problem: In January 2006 U.S. first-class mail rates increased to 39 cents for the first ounce, plus 24 cents for each additional ounce. If Sabrina spent $15.00 for a total of 45 stamps of these two denominations, how many stamps of each denomination did she buy?
-----------------
Quantity Equation: r + t = 45
Value Equation::::39r+24t = 1500 cents
------------------------------------
Multiply thru the 1st by 39 to get
39r + 39t = 39*45
---
Subtract that from the Value equation and solve for "t":
-15t = -255
t = 17 (# of twenty-four cent stamps)
-----------
Sinse r + t = 45, r = 45-17 = 28 (# of 39 cent stamps)
============================================================
Cheers,
Stan H.